OpenStudy (anonymous):

Find the area under 3x^2 = y from x=2 to x=5

OpenStudy (anonymous):

\[\int_{2}^{5} 3x^2 dx\]=\[(x^3)|_{2}^{5}\]=\[(5)^3-(2)^3\]=125-8=117